# Collective order boundedness of sets of operators between ordered vector spaces

### v3

@@ 2. $ii)$
Take any infinite dimensional Banach space $X$ with the trivial cone $\{0\}$, consider arbitrary discontinuous operator $T$ on $X$, and observe that $T$ is order to norm bounded simply because order intervals in $X$ are singletons.
@@ 3 Some applications to operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [14, p.75]).
@@
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
@@
Combining Propositions 3.1 and 2.1, and keeping in mind that every OBS with a generating normal cone is Archimedean, we obtain.
@@
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an OBS $X$ with a closed, generating, and normal cone. If $\mathcal{G}$ contains a (weakly) compact operator $S_{0}$ then $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Next, we consider commutative operator semigroups on AL-spaces.
@@
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$. Then $\|S\|\leq M<\infty$ for all $S\in{\text{\rm co}}(\mathcal{G})$. We conclude $\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{E}\subseteq MS_{0}B_{E}$. Since the set $S_{0}B_{E}$ is relatively weakly compact, it is almost order bounded by the Dunford-Pettis theorem (cf., [15][Theorem 2.5.4]). Thus, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. ∎
Combining Propositions 3.2 and 2.1 we obtain.
Corollary 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
We include the following convergence theorem.
Theorem 3.1.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an order continuous BL $E$. Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in\overline{{\text{\rm co}}}(\mathcal{G})$. Let $x\in E_{+}$. By the assumption, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded. Since $E$ is order continuous, $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is relatively weakly compact due to the Grothendieck lemma (cf., [2, Theorem 3.44]).
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [14, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
Corollary 3.3.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a one-parameter discrete operator semigroup on an order continuous BL $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs. The next two lemmas follow directly from Theorems 2.1, 2.2, and 2.3.
Lemma 3.1.
Every collectively order to norm bounded set of operators on an OBS with a closed generating cone is norm bounded.
Lemma 3.2.
Let $X$ be and OVS with generating cone. Then $\text{\bf L}_{rc}(X)=\text{\bf L}_{ob}(X)$. If $X$ is additionally Archimedean then $\text{\bf L}_{oc}(X)\subseteq\text{\bf L}_{ob}(X)$.
Next definition localizes several notions concerning semigrops.
Definition 3.1.
A semigroup $(T_{s})_{s\geq 0}$ on an OVS $X$ is
  1. $a)$
    locally order bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M$.
  2. $b)$
    locally o-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M$.
  3. $c)$
    locally ru-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{rc}(X)$ for every $M$.
A semigroup $(T_{s})_{s\geq 0}$ on a normed OVS $X$ is
  1. $e)$
    locally order to norm bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$
    for every $M$.
In several important cases, in each of the four items of Definition 3.1 the condition “for every $M$” can be replaced by “for some $M$”.
Proposition 3.3.
Let $(T_{s})_{s\geq 0}$ be a semigroup on an OVS $X$ such that $X_{+}-X_{+}=X$.
@@ 1. $i)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order bounded.
@@ 2. $ii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{rc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally ru-continuous.
@@ 3. $iii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order continuous.
@@ 4. $iv)$
    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order to norm bounded.
@@
Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order bounded.
$ii)$ By Lemma 3.2, it follows from $i)$.
@@
Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order continuous.
$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Let $n\varepsilon\geq M$ and take some $a\in X_{+}$. By Lemma 3.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$. Then
$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C\|a\|.$ (3)3( 3 )
Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order to norm bounded. ∎
In recent works [11, 12, 13] relatively uniformly continuous semigroups on VLs were deeply studied. The use of collectively qualified semigroups is too rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.

### v4

@@ 2. $ii)$
Consider an infinite dimensional Banach space $X$.
  1. $a)$
    Let $X_{+}:=\{0\}$ and take an arbitrary discontinuous operator $T$ on $X$. Then $T:(X,X_{+})\to X$ is order to norm bounded and order bounded simply because order intervals in $X$ are singletons. Clearly $X_{+}$ is closed but not generating.
  2. $b)$
    Pick a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of $X$, identify $X$ with $c_{00}(\Xi)$, and let $X_{+}:=\Big{\{}\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big{\}}$. Then $X_{+}-X_{+}=X$ but $X_{+}$ is not closed ((((e.g., by [7, Proposition 4.1])))). Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$. Clearly, $T$ is order to norm bounded and positive, yet not continuous.
Example 2.13.
Let $X$ be an ordered vector space with a generating cone which satisfies the dominated decomposition property ((((i.e., $[0,v_{1}]+[0,v_{2}]=[0,v_{1}+v_{2}]$ for every $v_{1},v_{2}$ in $X_{+}$)))), and let $Y$ be a Dedekind complete vector lattice. It is well known that the ordered vector space ${\cal L}_{ob}(X,Y)$ of order bounded operators of ${\cal L}(X,Y)$ is a Dedekind complete vector lattice (cf., [19, Theorem 83.4]). Let $A$ be an order bounded subset of ${\cal L}_{ob}(X,Y)$, say $A\subseteq[-T,T]$ for some positive operator $T\in{\cal L}_{ob}(X,Y)$. Let $x\in X_{+}$. Since $|Sz|\leq|S|z\leq Tz\leq Tx$ for all $z\in[0,x]$ and $S\in A$, then the set $A$ is collectively order bounded.
From the other hand, not every collectively order bounded set of operators is order bounded. Indeed, consider $X=\ell^{\infty}$, $Y=\mathbb{R}$, and $A=\{\delta^{n}:n\in\mathbb{N}\}\subset\ell^{\infty}_{+}$, where $\delta^{n}(\mathbb{x})=x_{n}$. The set $A$ is collectively order bounded, yet there is no $f\in\ell^{\infty}_{+}$ with $A\subseteq[0,f]$, since otherwise $f(1,1,\ldots,1,\ldots)\geq\sum_{k=1}^{n}\delta^{k}(1,1,\ldots,1,\ldots)=n$ for every $n\in\mathbb{N}$ that is absurd.
@@ 3 Some applications to operator semigroups
Under a semigroup $(T_{s})_{s\geq 0}$ on a vector space $X$ we understand a family of operators on $X$ satisfying $T_{s+t}=T_{s}T_{t}$ for all $s,t\in\mathbb{R}_{+}$, and $T_{0}=I_{X}$, where $I_{X}$ is the identity operator on $X$. We shall also use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on $X$, where $J=\mathbb{R}_{+},\mathbb{R},\mathbb{Z}_{+}$, or $\mathbb{Z}$. Such a semigroup is always commutative. More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [17, p.75]).
@@
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
@@
Theorem 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone, and assume that every order interval in $X$ is weakly compact. Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
Proof.
Since order intervals in $X$ are weakly compact then every almost order bounded subset of $X$ is relatively weakly compact due to the Grothendieck result (cf., [2, Theorem 3.44]). In particular, $\mathcal{G}$ be a collectively order to norm bounded. By Theorem 2.8, $\|S\|\leq M$ for some $M<\infty$ and all $S\in\mathcal{G}$.
Let $x\in X_{+}$. Then, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded, and hence relatively weakly compact. Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weak cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [17, Theorem 1.5]). Since $x\in X_{+}$ is arbitrary and $X_{+}$ is generating, $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
@@
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on an order continuous Banach lattice $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
The following example shows that the order continuity of the norm in $E$ is essential in Corollary 3.3.
Example 3.4.
(cf., [9, Exercise 2.2.22]) Consider the operator $T:c\to c$ such that $T(a_{1},a_{2},\ldots a_{k},\ldots)=\big{(}\frac{1}{2}a_{1},\frac{2}{3}a_{2},\ldots,\frac{k}{k+1}a_{k},\ldots\big{)}$. Clearly, $\big{\{}T^{n}\big{\}}_{n=1}^{\infty}\in\text{\bf L}_{ob}(c)$ for some $\varepsilon>0$. However, the Cesàro averages $\frac{1}{n}\sum_{j=0}^{n-1}T^{j}$ do not converge in the norm at $(1,1,\ldots 1,\ldots)\in c$.
The following corollary of Theorem 3.2 is immediate, because order intervals in the predual of a Neumann algebra are weakly compact by the Akemann theorem (cf., [9, Theorem 3.3.3]).
Corollary 3.5.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on the predual $\mathcal{M}_{*}$ of a von Neumann algebra $\mathcal{M}$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on ordered vector spaces.
Definition 3.6.
A semigroup $(T_{s})_{s\geq 0}$ on an ordered vector space $X$ is
  1. $a)$
    locally collectively order bounded ((((shortly, lco-bounded))))
    whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M>0$.
  2. $b)$
    lco-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M>0$.
  3. $c)$
    lcru-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ruc}(X)$ for every $M>0$.
A semigroup $(T_{s})_{s\geq 0}$ on a normed ordered vector space $X$ is
  1. $d)$
    lcon-bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$ for every $M>0$.
In several important cases, the condition “for every $M>0$” can be replaced by “for some $\varepsilon>0$” in each of the four items of Definition 3.6.
@@ 1. $d)$
Let ${\cal T}=(T_{s})_{s\geq 0}$ be a semigroup on an ordered vector space $X$ such that $X_{+}-X_{+}=X$.
@@ 1. $i)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-bounded.
@@ 2. $ii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ruc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcru-continuous.
@@ 3. $iii)$
    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-continuous.
@@ 4. $iv)$
    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcon-bounded.
@@
Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order bounded.
$ii)$ By Theorem 2.4, it follows from $i)$.
@@
Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order continuous.
$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$. Take $n\in\mathbb{N}$ with $n\varepsilon\geq M$. Let $a\in X_{+}$. By Theorem 2.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$. Then
$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C^{n}\|a\|.$ (3)3( 3 )
Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. As $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order to norm bounded. ∎
Under rather mild assumptions, every order bounded (o-continuous, ru-continuous, order to norm bounded) operator between ordered Banach spaces gives a rise to lco-bounded (resp., lco-bounded, lco-continuous, lcru-continuous, lcon-bounded) operator semigroup $(T_{t})_{t\geq 0}$.
Example 3.8.
Let $S:X\to Y$ be a bounded operator between Banach spaces. Let $Z=X\oplus Y$, define a bounded operator $A:Z\to Z$ by $A(x,y)=(0,Sx)$, and set $T_{t}=e^{tA}=\|\cdot\|$-$\sum_{k=0}^{\infty}\frac{t^{k}A^{k}}{k!}$ for $t\geq 0$. Then $(T_{t})_{t\geq 0}$ is a uniformly continuous semigroup on the Banach space $Z$ such that $T_{t}=I+tA$ for $t\geq 0$ because $A^{2}=0$.
Suppose $X$ and $Y$ are ordered Banach spaces and $X_{+}-X_{+}=X$.
  1. $(a)$
    Let $Y_{+}$ be normal and $S:X\to Y$ be order bounded. Then $S$ is order to norm bounded since $S$ is order bounded and $Y_{+}$ is normal. Theorem 2.8 implies that $S$, and hence $A$ is bounded. Since $A$ is order bounded, it follows from $T_{t}=I+tA$ for $t\geq 0$ that $(T_{t})_{t\geq 0}$ is lco-bounded.
  2. $(b)$
    Let $Y_{+}$ be normal and $S:X\to Y$ be ru-continuous. By Theorem 2.4, $S$ is order bounded, and hence $S$ is order to norm bounded because $Y_{+}$ is normal. Due to Theorem 2.8, $S$ is bounded. Thus $A$ is bounded. Since $A$ is ru-continuous, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcru-continuous.
  3. $(c)$
    Suppose additionally that $X$ is Archimedean. Let $Y_{+}$ be normal and $S:X\to Y$ be order continuous. By Theorem 2.1, $S$ is order bounded. Since $Y_{+}$ is normal, $S$ is order to norm bounded. Theorem 2.8 implies boundedness of $S$, and hence of $A$. Since $A$ is order continuous and $T_{t}=I+tA$, we have $(T_{t})_{t\geq 0}$ is lco-continuous.
  4. $(d)$
    Let $S:X\to Y$ be order to norm bounded. Then, by Theorem 2.8, $S$ is bounded, and so is $A$. Since $A$ is order to norm bounded, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcon-bounded.
In recent works of Glück, Kandić, Kaplin, and Kramar Fijavž [13, 15, 16] relatively uniformly continuous semigroups on vector lattices were deeply studied. Using of collectively qualified semigroups is too rough for getting progress in this direction. We postpone the investigation of relatively continuous semigroups on ordered vector spaces to a forthcoming paper.